OpenStudy (anonymous):

A particle is moving along a projectile path where the initial height is 96 feet with an initial speed of 16 feet per second. What is the maximum height of the particle?

OpenStudy (anonymous):

Assuming the initial speed is upwards, the position function is \[s(t)=-16t^2+16t+96\]

OpenStudy (anonymous):

To find the time of the maximum height, let t=-b/2a=-16/-32=1/2 sec

OpenStudy (anonymous):

Put that into the function to get the maximum height.

OpenStudy (anonymous):

wait so whats the answer

OpenStudy (anonymous):

Would you work out under his intruction?

OpenStudy (anonymous):

um ok...

OpenStudy (anonymous):

i just don't really get it

OpenStudy (anonymous):

We're right here for you :)

OpenStudy (anonymous):

Plug 1/2 into the function in the top post of the thread, and see what you get.

OpenStudy (anonymous):

t=1/2 s(t)=−16t2+16t+96 => s ( 1/2) = ...

OpenStudy (anonymous):

oh okay so s(1/2)= -16(1/2)^2 + 16(1/2) + 96 s(1/2) = 8^2 + 8 + 96 s(1/2)= 168 ?????

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

s (1/2) = 6 ( -t² + t + 6)

OpenStudy (anonymous):

where did 6 come from

OpenStudy (anonymous):

I just want to make it easier for you!

OpenStudy (anonymous):

Your order of operations is wrong. \[s(1/2) = -16(1/2)^2 + 16(1/2) + 96=-16(1/4)+8+96=-4+8+96=100\]

OpenStudy (anonymous):

I mean 16 :")

OpenStudy (anonymous):

I don't think six divides 16 evenly...

OpenStudy (anonymous):

OK.

OpenStudy (anonymous):

o ok so its 100?

OpenStudy (anonymous):

@nicole1217 all you do is plug the time in, the value S obtain is the height!

OpenStudy (anonymous):

Yep! high five :)

OpenStudy (anonymous):

yay! Thanks :)

OpenStudy (anonymous):

@nicole1217 Just want to emphasize the key idea that Ani. mentioned at first. "To find the time of the maximum height, let t=-b/2a=-16/-32=1/2 sec

OpenStudy (anonymous):

yeah i got it now